OpenStudy (anonymous):

hi, I'm having trouble with this problem. 4√8-√128+48+3√18

jimthompson5910 (jim_thompson5910):

What do you get when you simplify \[\large 4\sqrt{8}\]

OpenStudy (anonymous):

2 √2√4?

jimthompson5910 (jim_thompson5910):

what is the square root of 4

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

actually it should be \[\large 4\sqrt{8} = 4\sqrt{4*2}\] \[\large 4\sqrt{8} = 4\sqrt{4}*\sqrt{2}\] \[\large 4\sqrt{8} = 4*2*\sqrt{2}\] \[\large 4\sqrt{8} = 8\sqrt{2}\] Hopefully all that makes sense

jimthompson5910 (jim_thompson5910):

you would use this idea to simplify the other expressions

OpenStudy (anonymous):

still not quite understanding it.

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

i don't know how to start it, and i don't know the steps

jimthompson5910 (jim_thompson5910):

do you see how I simplified \[\large 4\sqrt{8} \]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

you would apply the same to the other pieces

jimthompson5910 (jim_thompson5910):

tell me what you get when you simplify \[\large \sqrt{128} \]

OpenStudy (anonymous):

8√16?

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

hint: 128 = 64*2

OpenStudy (anonymous):

2√8*8

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{128} = \sqrt{64*2}\] keep going

OpenStudy (anonymous):

what do i do with the 64?

OpenStudy (anonymous):

do i break it in half?

jimthompson5910 (jim_thompson5910):

then you use the idea that sqrt(x*y) = sqrt(x) * sqrt(y)

jimthompson5910 (jim_thompson5910):

so \[\large \sqrt{128} = \sqrt{64*2}\] \[\large \sqrt{128} = \sqrt{64}*\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

then what is the square root of 64?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

8*8√2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

the square root of 64 is 8, but you won't be left with 8*8

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{128} = \sqrt{64*2}\] \[\large \sqrt{128} = \sqrt{64}*\sqrt{2}\] \[\large \sqrt{128} = 8\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

see how I'm getting that

OpenStudy (anonymous):

slowly

jimthompson5910 (jim_thompson5910):

ok point out where you're stuck

OpenStudy (anonymous):

i don't know why you seperate the square roots in the second step

jimthompson5910 (jim_thompson5910):

because I'm using the rule \[\large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\] where x and y are nonnegative real numbers

jimthompson5910 (jim_thompson5910):

I'm using that rule to take advantage of the fact that the square root of 64 is 8

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

so the next one; is it: 2√48 = 2√6√2

OpenStudy (anonymous):

2*6√2

jimthompson5910 (jim_thompson5910):

list the perfect squares: 1, 4, 9, 16, 25, 36, 49, ... which perfect square (less than 48) is a factor of 48?

jimthompson5910 (jim_thompson5910):

what is the largest perfect square factor of 48

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

is that the only one? is that the largest?

OpenStudy (anonymous):

16 *3

OpenStudy (anonymous):

so 2√16*3

OpenStudy (anonymous):

2=√16√3

jimthompson5910 (jim_thompson5910):

square root of 16 is 4, so this means \[\large 2\sqrt{48} = 2\sqrt{16*3}\] \[\large 2\sqrt{48} = 2\sqrt{16}*\sqrt{3}\] \[\large 2\sqrt{48} = 2*4*\sqrt{3}\] \[\large 2\sqrt{48} = 8\sqrt{3}\]

OpenStudy (anonymous):

okay i'm getting there

OpenStudy (anonymous):

sorry i just woke up and i needed to get my brain into it

jimthompson5910 (jim_thompson5910):

now simplify \[\large 3\sqrt{18}\]

OpenStudy (anonymous):

3√9√9 3*3*3

OpenStudy (anonymous):

27?

jimthompson5910 (jim_thompson5910):

18 is NOT 9*9

OpenStudy (anonymous):

oops thats 81

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (anonymous):

3 √3*6

jimthompson5910 (jim_thompson5910):

is 3 a perfect square? is 6 a perfect square?

OpenStudy (anonymous):

3*6√3

jimthompson5910 (jim_thompson5910):

no you cannot do that

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

use 18 = 9*2

OpenStudy (anonymous):

3√9*2

jimthompson5910 (jim_thompson5910):

break up the root, then simplify the square root of 9 afterwards multiply

OpenStudy (anonymous):

3*3√2

OpenStudy (anonymous):

9√2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

So this means \[\large 4\sqrt{8} - \sqrt{128} + 2\sqrt{48} + 3\sqrt{18}\] turns into \[\large 8\sqrt{2} - 8\sqrt{2} + 8\sqrt{3} + 9\sqrt{2}\] now combine like terms

OpenStudy (anonymous):

do i combine the outside numbers?

jimthompson5910 (jim_thompson5910):

yes but only for the like terms

OpenStudy (anonymous):

so only the √2's

OpenStudy (anonymous):

8√3 9√2

OpenStudy (anonymous):

should there be a sign in between?

OpenStudy (anonymous):

8√3+9√2

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

you are 100% correct, so \[\large 4\sqrt{8} - \sqrt{128} + 2\sqrt{48} + 3\sqrt{18}\] simplifies to \[\large 8\sqrt{3} + 9\sqrt{2}\] In other words \[\large 4\sqrt{8} - \sqrt{128} + 2\sqrt{48} + 3\sqrt{18} = 8\sqrt{3} + 9\sqrt{2}\]

OpenStudy (anonymous):

holy cow!

jimthompson5910 (jim_thompson5910):

what's wrong?

OpenStudy (anonymous):

this was a radical nightmare

jimthompson5910 (jim_thompson5910):

well keep practicing and you'll get better

OpenStudy (anonymous):

But i'm glad you walked me through it, i needed to be able to visualize each step

OpenStudy (anonymous):

how do i give you a medal?

jimthompson5910 (jim_thompson5910):

by clicking the blue "best response" button

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

how do i distribute the square roots? i.e. √7(√6-√10)

jimthompson5910 (jim_thompson5910):

you multiply the outer term by each of the inner terms and simplify

jimthompson5910 (jim_thompson5910):

you are using the distributive property in your first step

OpenStudy (anonymous):

√63-√70 ? then simplify?

jimthompson5910 (jim_thompson5910):

7*6 is not 63